OpenStudy (anonymous):

the perimeter of a rectangle is 148 feet. the length is 14ft greater than one and a half times the width what is the length of the rectangle please help

OpenStudy (shelby1290):

the formula for perimeter of a rectangle is L + L + W + W = 2L + 2W = P

OpenStudy (shelby1290):

2L + 2W = 148 L is 14ft more than W x 1.5 OR L = 14 + 1.5W

OpenStudy (shelby1290):

L=?

OpenStudy (shelby1290):

2L + 2W = 148 2(14 + 1.5W) + 2W = 148

OpenStudy (shelby1290):

28 + 3W + 2W = 148

OpenStudy (shelby1290):

28 + 5W = 148 Subtract 28 from each side: 5W = 120 Divide both sides by 5: W = 24

OpenStudy (shelby1290):

to find L... L = 14 + 1.5 (W) L = 14 + 1.5 (24) L = 14 + 36 L = 50

OpenStudy (shelby1290):

check to make sure the left side =right side 2(50) + 2(24) = 148? 100 + 48 = 148? 148 = 148 its correct

OpenStudy (anonymous):

so the length of the rectangle is 50

OpenStudy (shelby1290):

yes

OpenStudy (anonymous):

thanks

OpenStudy (shelby1290):

no problem:)